# A Definition of AGI

### v2

@@ H.2 Visual Generation
“Generate an image with the following characteristics: Abraham Lincoln touches his toes while George Washington does chin-ups. Lincoln is barefoot. Washington is wearing boots.” [Marcus et al., 2022]
@@ H.3 Visual Reasoning
Please join the circles together to form a letter (ignore the squares). [Brown et al., 2019]
@@
Answer: A [Maeda et al., 2013]
@@
Answer: Blue [Gemini Robotics Team et al., 2025]
@@
Answer: CC daily 4.00PM [Wang et al., 2024]
@@
SPACE [Ramakrishnan et al., 2025], a test of various visual reasoning skills, must be above 80%.
SpatialViz-Bench [Wang et al., 2025], a test of mental rotation and folding, must be above 80%.
CharXiv [Wang et al., 2024], a test of figure question and answering, must be above 80%.
ERQA [Gemini Robotics Team et al., 2025], a test of embodied reasoning, must be above 80%.
ClockBench [Safar, 2025], a test of reading clock hands, must be above 80%.
@@ I.1 Phonetic Coding
Repeat the following nonsense word: C phontetic_coding.wav
@@ I.2 Speech Recognition
Transcribe this TED talk: U “Can we build Al without losing control over it?”

### v3

@@ H.2 Visual Generation
“Generate an image with the following characteristics: Abraham Lincoln touches his toes while George Washington does chin-ups. Lincoln is barefoot. Washington is wearing boots.” [dalle2_failures]
@@ H.3 Visual Reasoning
Please join the circles together to form a letter (ignore the squares). [brown2019test]
@@
Answer: A [spatial_reasoning_ijee]
@@
Answer: Blue [gemini_robotics]
@@
Answer: CC daily 4.00PM [charxiv]
@@
SPACE [ramakrishnan2025doesspatialcognitionemerge], a test of various visual reasoning skills, must be above 80%.
SpatialViz-Bench [spatialviz], a test of mental rotation and folding, must be above 80%.
CharXiv [charxiv], a test of figure question and answering, must be above 80%.
ERQA [gemini_robotics], a test of embodied reasoning, must be above 80%.
ClockBench [Safar_ClockBench_2025], a test of reading clock hands, must be above 80%.
@@ I.1 Phonetic Coding
Repeat the following nonsense word: C phonetic_coding.wav
@@ I.2 Speech Recognition
Transcribe this TED talk: U “Can we build AI without losing control over it?”
@@
LibriSpeech [librispeech] (test-clean) WER must be less than 5.83% (human-level) for the clean audio 2%.
LibriSpeech [librispeech] (test-other) WER must be less than 12.69% (human-level) for the noisy audio additional 2%.